OpenStudy (anonymous):

Find the area between the curves on the given interval. y = x3 – 1, y = 1 – x, -2 ≤ x ≤ 2

OpenStudy (anonymous):

this is going to suck because you have to divide the region up in to several intervals

OpenStudy (anonymous):

wow okay

OpenStudy (anonymous):

you can see that they switch places, so you are going to have to compute \[\int_{-2}^{-1}(x-1)-(x^3-1)dx\] \[\int_{-1}^0(x^3-1)-(x-1)dx\] \[\int_0^1(x-1)-(x^3-1)dx\] \[\int _1^2(x^3-1)-(x-1)dx\]

OpenStudy (anonymous):

each of these is very easy, but doing 4 of them is a drag

OpenStudy (anonymous):

so i will integrate them or?

OpenStudy (anonymous):

yes, but of course simplify them first

OpenStudy (anonymous):

each is either \(x^3-x\) or \(x-x^3\) depending

OpenStudy (anonymous):

so the anti derivative will either be \(\frac{x^4}{4}-\frac{x^2}{2}\) or the other way around it is easy, just a pain

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

ok so i got -x^3 - 3/2 -x - 1/4 1/2 - x^3 and 15/4 - x